1.   Calculate the [OH-] and pH of each of the following strong base solutions:

 a.   0.0050 M KOH

b.   2.055 g of KOH in 500.0 mL of solution

c.    1.00 mL of 0.175 M NaOH diluted to 2.00 L.

 


1
Expert's answer
2021-10-22T03:50:42-0400

We are given 0.0050 M KOH solution:


KOH is a strong base, so it dissociates in the solution 100% as follows:


KOH→K++OH-


The concentration of hydroxide ion is equal to the that of KOH because 1 mole of KOH releases 1 mole of hydroxide ion.


[OH-]=0.0050M

pOH= -log(OH-)

= -log0.005

pOH= 2.3

pH= 14-2.3= 11.3